# ANGO: A Next-Level Evaluation Benchmark For Generation-Oriented Language Models In Chinese Domain

### 10 Jan 2024 v1

@@ 2.1 Data Collection
After collecting without any manipulation, we obtained a total of 674,675 records. In order to enhance the quality of our data, we employed a series of simple but efficient preprocess functions, as shown in Figure 2: (1) Mock exams often include previous exam questions, which lead to numerous duplicates. We remove duplicates based on the content of records; (2) Since our primary focus was on NLP instead of Multi-Modal, we removed all records containing pictures in both the questions and options; (3) By manual recognition aided by OCR model, we extracted 5,574 unique formulas from a pool of 34,062 formula images.
@@ 2.2.1 Keypoint Tree
The keypoint categorization can comprehensively represent the distribution of capabilities required to complete a question. 5 major categories were identified in ANGO as the first-level keypoints.
@@ 2.2.2 Difficulty Measurement
With this standard, we categorized the difficulty scores into nine distinct levels, as shown in Figure 4, ranging from Level 1 (easiest) to Level 9 (most challenging).
@@ 3.1 Few-Shot History Sample
Due to the distinct category hierarchy in our keypoint tree architecture and the insufficient number of records for some $K$, we design a dynamic length example sampling strategy that ensures minimal information loss and divides into below 4 steps. The next step will only be taken if there is not any available record in current step.
@@
  • Step 3-$K^{p}$: Iterate through the parent nodes $K^{p}$ within $P(K)$ and repeat Step 2 on each $K^{p}$.
  • Step 4-$K^{g}$: Reiterate through the parent nodes $K^{p}$ within $P(K)$, repeat Step 2 on each $K^{g}$ of $G(K^{p})$.
Benefits from the multi-step sampling approach, it ensures that the examples in the beginning are more closely aligned with the test record compared to those at the end.
@@ 3.2 Testset Sample
Since $K_{s}$ may contain more than one $K$ and the substantial overlap across different $K_{s}$, sampling records for each $K$ or $K_{s}$ will definitely lead to a bad result. To achieve a more balanced distribution, we propose a sample strategy in following steps:
@@
After calculation, select the $K_{s}$ with the minimum distance, randomly choose a record, and insert it into the test. Concurrently, update $V_{C}$ by adding the selected $V_{K}$. Just repeat this step until each $K$ has $N$ records or there are no more records available for $K$.
@@ 4.1 Metric
We choose accuracy to be the ultimate metric for assessing the performance of our models.
For the base model(Pre-Trained stage), as there are questions with multiple correct options in our dataset, it’s not possible for us to calculate the logits of different options to get the answer, So we simply split the generated output of the model by newline delimiters and extracted the first element as the answer to match the correct answer.
For chat models(Supervised Fine Tune stage), models often failed to provide answers in our expected format. To address this issue, we split the generation output based on Chinese and English tags and designed some regular expressions to capture options with a special format.
@@
Evaluating models solely based on their ability to provide correct answers does not capture the full spectrum of human-like behavior and thinking patterns. While accuracy is important, overlooking the value in identifying common human errors may hinder progress toward building truly human-centric AI. This motivated us to propose new evaluation metrics to address this limitation.
@@ 4.2 Prompt
Similar to the settings of other benchmarks, ANGO adopts the few-shot as an evaluation strategy with history examples sampled in Section 3.1. By exposing the model to a limited number of examples during the evaluation process, we can effectively gauge its generalization and adaptability to novel scenarios. Furthermore, the few-shot evaluation strategy offers practical advantages when it comes to extracting the model’s output options, an example of a prompt is shown in Figure 5.
@@
In particular, we have made a deliberate decision to minimize the inclusion of Chain Of Thought(CoT)[15] in our evaluation procedures, with the exception of the API over which we lack control. This decision stems primarily from prior research[10] that has unequivocally demonstrated the detrimental impact of CoT on model performance.
In cases where the length exceeds this limit, we systematically reduce the number of examples from the end to the beginning. Thanks to our example sampling strategy(Section 3), the examples at the head provide more informative content compared to those at the tail.
@@ 4.3 Framework
  • Seasonal Dynamic Evaluation: Thanks to our optimized sampling strategy, we can periodically refresh a portion of the testset and update the leaderboard so that each season has a distinct testset to prevent data leaks.
@@
  • Question Elimination: By calculating questions’ accuracy across models and seasons. Questions exceeding 100% accuracy are temporarily removed to ensure sufficient discriminability.
@@ 5.2 Models
In light of limited computational resources, we were compelled to utilize a single A100 unit for conducting our experiments. As a result, we were obliged to select the following models of relatively smaller scale or models accessible through API services.
@@ A.1 Expression&Comprehension
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句排序题 Statement Ordering Questions 非首句特征 First Sentence Features
@@ A.2 Quantitative Relations
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 排列组合问题 Permutation And Combination Problems 不相邻问题 Adjacent Issues

### 21 Feb 2024 v2

@@ 2.1 Data Collection
After collecting without any manipulation, we obtained a total of 674,675 records. In order to enhance the quality of our data, we employed a series of simple but efficient preprocess functions, as shown in Figure 2: (1) Mock exams often include previous existed exam questions, which lead to numerous duplicates. We remove duplicates based on the content of records; (2) Since our primary focus is on NLP instead of Multi-Modal, we removed all records containing pictures in both the questions and options; (3) By manual recognition aided by OCR model, we extracted 5,574 unique formulas from a pool of 34,062 formula images and convert them to text.
@@ 2.2.1 Keypoint Tree
The keypoint categorization can comprehensively represent the distribution of capabilities required to complete a question. These 5 major categories are identified in ANGO as the first-level keypoints.
@@ 2.2.2 Difficulty Measurement
With this standard, we divided questions in ANGO into 9 distinct levels, as shown in Figure 4, ranging from Level 1 (easiest) to Level 9 (most challenging).
@@ 3.1 Few-Shot History Sample
Given the unique category hierarchy in our keypoint tree architecture, coupled with a limited number of records for certain questions’ $K_{s}$ , we designed a dynamic length example sampling strategy. This strategy ensures minimal information loss and is split into four steps. Proceeding to the next step only occurs if no available records exist in the current step.
@@
  • Step 3-$K^{p}$: Iterate through the $K$ within $K_{s}$ as a loop and sample one record associated with each $K^{p}$ in $P(K)$.
  • Step 4-$K^{g}$: Iterate through the $K$ within $K_{s}$ as a loop and sample one record associated with each $K^{g}$ in $G(K)$.
Benefits from the multi-step sampling approach, it ensures that the examples in the beginning are more closely aligned with the target question compared to those at the end.
@@ 3.2 Testset Sample
Since $K_{s}$ may contain more than one $K$ and and there’s significant overlap among different $K_{s}$, sampling records for each $K$ or $K_{s}$ will definitely lead to a imbalance result. To obtain a more even distribution, we suggest the following strategic sampling method:
@@
After calculation, select the $K_{s}$ with the minimum distance, choose a record with the difficulty level that appears the least frequently in the current testset, and insert it into the test. Concurrently, update $V_{C}$ by adding the selected $V_{K}$.
Just repeat Insert and Update until each $K$ has $N$ records or there are no more records available for $K$.
@@ 4.1 Metric
We choose accuracy to be the metric for assessing the performance of our models.
For the base model(Pre-Trained stage), as there are questions with multiple correct options in our dataset, it’s not feasible for us to compute the logits of varying options to derive the answer, we opted to divide the model’s generated output by newline delimiters, subsequently extracting the first element as the model’s answer.
For chat models(Supervised Fine Tune stage), models often failed to provide response in our expected format. To rectify this problem, we split the generated output based on various regular expressions that include Chinese and English tags, then manually vetted and captured the options as the model’s answer.
@@
Evaluating models solely based on their ability to provide correct answers does not reflect the full spectrum of human-like behavior and thinking patterns. While accuracy is important, overlooking the value in identifying common human errors may hinder progress toward building truly human-centric AI. This motivated us to propose a new evaluation metric to address this limitation.
@@ 4.2 Prompt
Similar to the settings of other benchmarks, ANGO adopts the few-shot as an evaluation strategy with history examples sampled in Section 3.1. By exposing the model to a limited number of examples during the evaluation process, we can effectively gauge its generalization and adaptability to novel scenarios. Furthermore, the few-shot evaluation strategy offers practical advantages when it comes to extracting the model’s output options. Prompt example is shown in Figure 5.
@@
We have consciously chosen to minimize the incorporation of the Chain Of Thought (CoT)[15] in our evaluation processes, excluding the API which remains beyond our control. This decision mainly stems from previous research[10] which conclusively evidenced the negative impact of CoT on model performance.
When the length surpasses the limit of model, we methodically reduce the quantity of examples from the end towards the beginning. Thanks to our sample strategy (outlined in Section 3), the examples at the begin offer more informative content when compared to those placed at the end.
@@ 4.3 Framework
  • Seasonal Dynamic Evaluation: We will periodically refresh a portion of the testset and update the leaderboard so that each season has a distinct testset to prevent data leaks.
@@
  • Question Elimination: By calculating questions’ accuracy across models and seasons. Questions exceeding 100% accuracy will temporarily be removed to ensure sufficient discriminability.
@@ 5.2 Models
In light of limited computational resources, we were compelled to utilize a single A100 unit for conducting our experiments. As a result, we were obliged to select the following models as our candidates.
@@ A.1 Expression&Comprehension
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句排序题 Statement Ordering Questions 非首句特征 Non-First Sentence Features
@@ A.2 Quantitative Relations
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 排列组合问题 Permutation And Combination Problems 不相邻问题 Non-Adjacent Issues